OpenStudy (shelld93):

Need confirmation of an answer! the question was, rearrange to make y the subject d = 7y - 4 I ended up with d = 7y - 4 d + 4 = 7y - 4 + 4 d + 4/7 = 7y/7 d + 4/7 = y y = d+4/7

OpenStudy (anonymous):

y=(d+4)/7

OpenStudy (anonymous):

up to here its correct d + 4 = 7y - 4 + 4

OpenStudy (shelld93):

are the brackets needed ? How will I know if they are or not ? What did I do wrong ?

OpenStudy (anonymous):

but for the line d + 4/7 = 7y/7 correct form is (d+4)/7 = 7y/7

OpenStudy (anonymous):

yup the brackets are neede when dividing or multiplying this way..

OpenStudy (shelld93):

okay so brackets for division and multiplication and i dont understand your other reply.

OpenStudy (shelld93):

it looks like text is missing

OpenStudy (anonymous):

hope its easy for you now..

OpenStudy (anonymous):

which statement..

OpenStudy (shelld93):

but for the line d + 4/7 = 7y/7 correct form is (d+4)/7 = 7y/7

OpenStudy (imstuck):

The answer is d + 4/7 = y.

OpenStudy (anonymous):

i meant to say that d + 4/7 = 7y/7 is incorrect and (d+4)/7 = 7y/7 is correct form

OpenStudy (shelld93):

okay so i went wrong because i didnt have brackets ? thats all that is wrong with my answer ?

OpenStudy (perl):

looks like, use parentheses when in doubt :P

OpenStudy (anonymous):

@ShellD93 yup

OpenStudy (shelld93):

thank you :)